### v1

# Puzzle Similarity: A Perceptually-guided No-Reference Metric for Artifact Detection in 3D Scene Reconstructions

Abstract
Modern reconstruction techniques can effectively model complex 3D scenes from sparse 2D views. However, automatically assessing the quality of novel views and identifying artifacts is challenging due to the lack of ground truth images and the limitations of no-reference image metrics in predicting detailed artifact maps. The absence of such quality metrics hinders accurate predictions of the quality of generated views and limits the adoption of post-processing techniques, such as inpainting, to enhance reconstruction quality. In this work, we propose a new no-reference metric, Puzzle Similarity, which is designed to localize artifacts in novel views. Our approach utilizes image patch statistics from the input views to establish a scene-specific distribution that is later used to identify poorly reconstructed regions in the novel views. We test and evaluate our method in the context of 3D reconstruction; to this end, we collected a novel dataset of human quality assessment in unseen reconstructed views. Through this dataset, we demonstrate that our method can not only successfully localize artifacts in novel views, correlating with human assessment, but do so without direct references. Surprisingly, our metric outperforms both no-reference metrics and popular full-reference image metrics. We can leverage our new metric to enhance applications like automatic image restoration, guided acquisition, or 3D reconstruction from sparse inputs.
1 Introduction
Image-based rendering and 3D reconstruction from a sparse set of 2D views has received ample attention in recent years, both for pure geometry reconstruction and radiance-field modeling. Classical approaches using simple triangulation and epipolar geometry through methods such as structure from motion (SfM) to produce sparse point-clouds of diffuse color [28]. Densifying these representations can be done explicitly [13]. Alternatively, one can learn continuous, implicit representations [22, 3, 25], normally modeled through some kind of multi-layer perceptron (MLP). A tangential problem to these efforts is the collection of 2D data, and the handling of corrupted, distorted, or simply incomplete sets of images from an object or scene we would like to model. Learning representations from very sparse inputs has been a widely studied topic as well [44, 4, 38, 5]; they normally leverage learned priors on large datasets, helping to fill in the gaps by enforcing 3D consistency and that resulting reconstructions follow natural image statistics. However, quantifying the quality of novel views from these reconstructions is still problematic. These views can contain artifacts due to the sparsity of the dataset their model was constructed from, and automatically identifying them helps with restoration (e.g. masking for image-based inpainters [32]) or simply to guide future data acquisition to fill the gaps [15]. Recent works have followed a bayesian approach to quantification of the uncertainty of an area belonging or not to a reconstructed model or scene [7], which could potentially be leveraged for simple artifact detection; however, they require implicit representations of the scene, with fundamental changes to the scene model, and are not practical for more general applications that require visual artifact identification outside of scene reconstruction, as well as incapable of detecting artifacts not arising from lack of coverage.
To tackle this, we propose a novel approach for artifact detection that can be leveraged on any set of images without an encoded explicit or implicit model of the scene or object they depict. As opposed to visual difference predictors (VDPs) [18] (which require references) and no-reference quality metrics [23, 24] (which typically do not provide maps, but rather produce single values of overall quality) our approach provides visual artifact maps with no direct references. We leverage learned perceptual patch statistics from small clean datasets and compare them to the embdedded statistics of new images from a similar distribution (i.e. novel reconstructed views from a 3DGS [13] representation) to obtain artifact maps without direct references. We test our generated maps through a human experiment where we ask participants to manually identify artifacts and distortions in images to generate ground-truth data of visual artifacts. Our results show that our method agrees with human assessment, correlating better than both no-reference and full-reference metrics. To summarize, our contributions are the following:
  • A novel no-reference visual quality/artifact identification metric particularly tailored for image-based rendering,
  • a novel dataset of human-assessed artifact and distortion identification to validate our metric,
  • and an application on image restoration and 3D reconstruction enhancement that showcases the utility of our approach.
2 Related Work
3D reconstruction and Image-based rendering
Reconstructing 3D objects or scenes from sparse sets of 2D observations is a fundamental problem in vision [18]. Particularly, in the context of novel view synthesis, the objective is to approximate the radiance field (i.e. 5D function encoding spatially varying radiance emission) of specific objects or scenes. Most methods however, differ either on the model used to encode the function, or the rendering procedure. Implicit approaches model the radiance field as a continuous function, approximated by a multi-layered perceptron (MLP) [22, 31]. Rendering is usually done via sampling the implicit volume using ray-marching [36], which provides spatially varying values of density and anisotropic color emission modeled through Spherical Harmonics. Improvements over this formula have tackled performance limitations, either by using more efficient sampling techniques [27, 8, 25] or by distilling the implicit space into explicit density and anisotropic appearance volumes [43, 39]. On the other hand, purely explicit models do not require any pre-training using implicit functions, and were originally Eulerian in nature [42]. Explicit models are easier to optimize, usually faster, and more interpretable, which can help in different tasks such as scene editing or animation. More recently, anisotropic Lagrangian approaches have found tremendous success [13]. However, these explicit methods have introduced some limitations of their own along the way. Methods like 3D Gaussian Splatting [13] can only model areas that are directly supervised, and degrade less gracefully than implicit counterparts when querying viewpoints substantially outside the training set coverage. Detecting artifacts arising from the lack of coverage is difficult due to the lack of reference images. Our method produces these masks via supervision on the training data solely, which can enable unsupervised restoration (automatic inpainting of the artifacts based on available context [6]) or simply automatically guide further image acquisition to complete the dataset efficiently [15].
Image Quality Metrics (IQMs)
We can roughly categorize IQMs on their reliance on reference images, or the lack of it. Full-reference techniques traditionally input a reference image and its distorted counterpart and compute pixel or patch wise differences, which can either be pooled across an image to provide a single quality value [37, 49] or in some cases, kept as a pixel-wise map to visualize the location or source of the distortions. Classic examples also include mean-absolute error (MAE) or mean-squared error (MSE). Pixel-map metrics share similarities with visual difference predictors (VDPs), with the distinction usually being that VDPs leverage some explicit or implicit model of the human visual system (HVS) to align distortion predictions with human assessment. In some cases, it is also possible to transform usually pooled metrics like SSIM [37], FSIM [46] and LPIPS [49] into pixel maps, and we show how in Section 4.1.
No-reference visual quality metrics relinquish the necessity of direct references, and are most commonly learned, often relying on mean-opinion based supervised learning. Alternatively to leveraging datasets on human quality assessment [24, 12], they can also be supervised on extracted features from natural image statistics [23, 40], or even synthetic scores [41]. They are not usually capable of producing spatial quality maps, which makes it impossible to determine the source of reduced quality or localize the potential artifacts.
Some IQMs however, are capable of producing visual maps as well [26, 40, 12]; this makes them the closest to our work. The most relevant work in this space, PIQE [26] similarly attempts to quantify visual artifacts without relying on any supervision; rather, it extracts local features from image patches and quantifies quality as a measure of distortions in the patch, in accordance with a low-level human vision model. In contrast, our work does not make any assumptions on the HVS, rather leveraging the latent space of a model pretrained on natural images to measure the euclidean distance in feature space of candidate image patches to a limited set of image patches from a similar distribution (i.e., images from the same scene in the context of scene reconstruction). This enables us to obtain a higher level of alignment with human assessment.
Visual Difference Predictors (VDPs)
VDPs usually differ from IQMs in the more explicit integration of models of human vision, leveraging the naturally compressive nature of the human visual system [18] to predict perceived differences between images. Improvements over the original framework have increased its robustness by extending their applicability to high dynamic range imagery [20]; making them eccentricity and motion-aware [19]; and integrating perceived color [21]. VDP frameworks have seen use in foveated rendering applications [35] and perceptually-aware tone mappers [34]. In contrast to VDPs, our model relinquishes the necessity of direct references, while performing similarly to state of the art VDPs such as FovVideoVDP [19].
3 Our Method
Figure 1: Schematic representation of our Puzzle Similarity metric.
Let us establish an analogy for our method: pretend each reference image is a puzzle with many puzzle pieces. To test if a new image is similar to our references, we would simply shuffle all pieces from all puzzles from our references and try to reassemble the test image only using those pieces. If the new image is very similar to the references, we should have enough puzzle pieces to compose the other image confidently. However, if the image holds regions very different from what we saw in the reference images, we would lack puzzle pieces to assemble this area, effectively leaving holes in the newly assembled puzzle (image). An overview of our approach through this analogy can be seen in Figure 1.
In our work, the puzzle pieces correspond to embedded image patches. In order to assess patch similarity, an obvious approach would involve computing the dot product between all patches; best-matching pieces would be recorded to create a similarity map. This simplistic approach, however, would hardly align with human assessment. Inspired by the close correlation between human quality judgment and latent CNN feature maps [33, 49], we employ a pre-trained convolutional network [16, 10, 30] to embed all the references, computing similarity in the latent feature space. Note that comparing feature map ”pixels” in a convolutional network is similar to comparing individual patches in the input domain; this is due to the locality of the sliding kernels when convolving. The patch size is dependent on the receptive field (showcased in Figure 2).
Figure 2: Receptive field of a multi-layer CNN. Note how one pixel in the last layer can be seen as an embedded patch of the input space.
Choice of layers
Choosing the right layers for embedding is essential to maximize the quality of the predicted spatial maps. While early layers feature small receptive fields and capture fine details, deeper layers have larger receptive fields and capture coarser features. This can be observed in Figure 3, where we showcase different VGG layers. It is essentially a trade-off between prediction granularity, accuracy and speed. We identified that combining multiple layers into our metric computation incorporates the various levels of abstraction and scales in a robust manner. We thus compute the weighted average of the three layers; we empirically found that halving the image resolution more than three times did not significantly improve our results as the scale becomes too small and the pool of reference vectors too little and specific to find good correspondences among novel images, even for well-reconstructed areas. This observation suggested that features from the layers before the third down-sampling were most useful for our cause.
(a) Rendering
(b) $2^{\text{nd}}$ VGG layer
(c) $3^{\text{rd}}$ VGG layer
(d) $4^{\text{th}}$ VGG layer
Figure 3: Puzzle Similarity computed only on a single VGG layer respectively. Note how the second layer has a much better resolution and generally colder colors, while the fourth layer is much smoother and features a wider range of values. Warm colors indicate the presence of artifacts or poor reconstruction quality.
Computing patch similarity
To compute the similarity map of a test image, we feed all references and the test image through a pre-trained network $\mathcal{F}$ to obtain the embeddings. We repeat the exact computation for each network layer, so we will describe the steps once for one layer $\ell$. To find the similarity $s_{\ell}(x,y)$ of the best matching puzzle piece for a pixel of the embedded test image at some spatial location $(x,y)$, we compute the cosine similarity based on the feature vector $\mathcal{F}_{\ell}(x,y)$ and all other feature vectors of all $N$ reference images of the same layer $\ell$ and select the correspondence with the highest similarity:
$s_{\ell}(x,y)=\max_{n,x^{\prime},y^{\prime}}\hat{\mathcal{F}}_{\ell}(x,y)\cdot\hat{\mathcal{F}}^{(n)}_{\ell}(x^{\prime},y^{\prime})$ (1)
where $\hat{\mathcal{F}}$ denotes a feature vectors scaled to unit length ^$\hat{\mathcal{F}}_{\ell}(x,y)=\frac{\mathcal{F}_{\ell}(x,y)}{||\mathcal{F}_{\ell}(x,y)||_{2}}\in\mathbb{R}^{C_{\ell}}$ and $\cdot$ is the dot product. Note that we compute the cosine similarity with any feature vector of the same layer from all references, not just those at the same spatial position. This relinquishes spatial relations and makes the method robust to simple camera movements that only shift the image horizontally or vertically. We iterate this maximum search for all pixels of the test image’s embedding to construct the similarity mask $S_{\ell}$.
$S_{\ell}(\mathcal{I})=\begin{bmatrix}s_{\ell}(1,1)&s_{\ell}(1,2)&\cdots&s_{\ell}(1,W_{\ell})\\ s_{\ell}(2,1)&s_{\ell}(2,2)&\cdots&s_{\ell}(2,W_{\ell})\\ \vdots&\vdots&\ddots&\vdots\\ s_{\ell}(H_{\ell},1)&s_{\ell}(H_{\ell},2)&\cdots&s_{\ell}(H_{\ell},W_{\ell})\\ \end{bmatrix}$ (2)
where $\mathcal{I}$ is the test image. We repeat this for a set of layers and combine them into a final quality map. To match the spatial dimensions of each layer, we bilinearly upsample each map to the original image size and combine them with an affine combination:
$S(\mathcal{I})$ $=\sum_{\ell}w_{\ell}\;\text{Upsample}\left(S_{\ell}(\mathcal{I},\;\mathcal{I}_{\text{ref}}^{1:N})\right)$ (3)
with $\sum_{\ell}w_{\ell}=1$ and reference images $\mathcal{I}_{\text{ref}}^{1:N}$. To utilize optimized hardware, please note how the computation of $S_{\ell}$ can also be expressed as an outer product between the spatially flattened embeddings:
^$\hat{\mathcal{F}}_{\ell}(\mathcal{I}^{1:N})$ $\in\mathbb{R}^{N\times H_{\ell}\times W_{\ell}\times C_{\ell}}$ (4)
~$\tilde{\mathcal{F}}_{\ell}(\mathcal{I}^{1:N})$ $=\text{flatten}\left(\hat{\mathcal{F}}_{\ell}(\mathcal{I}^{1:N})\right)\in\mathbb{R}^{NH_{\ell}W_{\ell}\times C_{\ell}}$
[…]
where the test image $\mathcal{I}$ is a special case with $N=1$, $\otimes$ is the outer product, and rowmax applies the max over the first dimension. While a naïve implementation of this outer product would require substantial amounts of memory for larger $N,H,W$, we provide an efficient implementation through blockwise tiling with intermediate max-reduction, which we detail in the Supplemental.
Pre-trained Model Choice
The choice of pre-trained neural network, through which the embeddings will be created, is a key component of our work. We primarily considered classic models including VGG-16, VGG-19, AlexNet, and SqueezeNet [30, 16, 10]. Some of the critical considerations are model complexity and memory requirements, which we summarized in the Supplemental, as well as their specific tested performance in our human assessment alignment task. Beyond quality performance, reducing the memory footprint and computational complexity is key as it may impact the possibility of downstream applications of our metric, which, given its differentiability, could be leveraged in optimization procedures.
We empirically found that while VGG produces the most fine-grained maps, AlexNet and SqueezeNet still managed to perform similarly, with the key advantage of doing so at a substantially reduced computational cost. We opted for SqueezeNet as it aligned best with our test examples, specifically using layers $\ell\in\{2,3,4\}$ with the weights $w_{2}=0.67$, $w_{3}=0.2$, and $w_{4}=0.13$, which we found heuristically.
4 Results
We will now analyze how our method stacks against competing approaches for both full-reference and no-reference visual map prediction in the context of reconstruction and image-based rendering.
In order to quantify the correlation between all these maps and human assessment, we present a novel dataset on human artifact identification, which we manually collected and can be found here 111www.placeholder__\__fake__\__link__\__for__\__dataset.comto facilitate future research on the topic.
As for our method, for each different scene, we compute embeddings on their respective training dataset and leverage each of them respectively to compute visual quality maps on validation and artifact-ridden views, as explained in Section 3.
4.1 A Novel Artifact Identification Dataset
We created a dataset of human-perceived artifacts in 3D reconstructed views with corresponding ground truths collected through a user study. In order to generate images with typical reconstruction artifacts, we run 3D Gaussian Splatting [13] on twelve scenes from Mip-NeRF360 [3], Tank and Temples [14] and Deep Blending [9] datasets, using default parameters but withholding substantial amounts of training images. Omitting training views increases the chance of artifacts appearing in those withheld views, while still allowing us to have clean references. For each dataset, we selected three renderings that demonstrated a mix of well-reconstructed areas, strong artifacts, and subtle artifacts, resulting in 36 samples across 12 datasets.
Experiment details
We asked 23 participants to segment visible artifacts in each of the 36 sample images under controlled viewing conditions using the tool developed by Wolski et al. [1], which the authors kindly provided. We include details on the participants’ self-reported gender and age distributions in Supplementary material, as well as detailed viewing and display conditions. During the experiment, users had no undistorted, artifact-free references at their disposal and thus had to judge individual images at face value. They would then mark areas found to be unnatural or unappealing, creating a binary mask.
With the dataset, we can evaluate the agreement between human judgment and any metric output by simply averaging all binary masks to estimate the probability of each pixel being marked as an artifact. Figure 4 shows example renderings (a) alongside metric predictions (b)-(d) and their average human-produced mask (e).
We evaluate our method against both full-reference and no-reference IQMs, and the state-of-the-art VDP. To assess their alignment with human perception, we will correlate their maps with the human ratings from our dataset, which we described in Section 4.1.
4.2 Comparison with Full-Reference Metrics
Adapting pooled metrics to produce spatial quality maps
In terms of single-valued quality metrics, we selected commonly used reconstruction quality metrics (L1, L2, SSIM [37], LPIPS [49]). We adapt L1, L2 and SSIM by simply removing the pooling step. For LPIPS, after computing the distance in the embedding space, the metric already upsamples each feature map back to the original image resolution. We can then pool across maps but not along 2D image dimensions, allowing us to retain the map. We compare LPIPS with three base models: VGG, AlexNet, and SqueezeNet. We further compare with F LIP [2], a commonly found metric in quality assessment, particularly in the field of physically-based rendering; it already produces spatial quality maps by default. The no-reference metric CNNIQA [12] was applied on patches as described in their paper. We applied padding to avoid cropping the borders and upsampled the final map. PIQE [26] already produces three different kinds of maps that we averaged.
To correlate them to our human ratings, we first compute each metric map for each sample rendering and then compute the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC). To account for the different domains of the compared metrics and possibly non-linear relations, we fit a 5-parameter logistic curve for a fair comparison as suggested by [1, 17, 47, 29]:
$q(x)=a_{1}\left\{\frac{1}{2}-\frac{1}{1+\exp{(a_{2}(x-a_{3}))}}\right\}a_{4}x+a_{5}$ (5)
where $x$ is an individual quality score and $a_{1\dots 5}$ are the tunable parameters optimized using gradient ascent to maximize either PCC or SRCC.
(a) Artifact-ridden rendering
(b) PIQE map
(c) FovVideoVDP map
(d) PuzzleSim map (ours)
(e) Human labeling
Figure 4: Selection of image quality maps for artifact-ridden renderings from various scenes. The last column shows ground-truth human assessments from our collected dataset.
Results
Puzzle Similarity consistently achieves high correlation with human-perceived artifacts across all datasets, outperforming well-established full-reference (FR) metrics such as SSIM, L1, and L2 norms, even without access to direct ground-truth correspondences. Tables 1 & 2 report the average Pearson and Spearman correlations for each scene. Our dataset includes three images per scene, with artifact types varying per scene. For example, the garden scene has prominent black regions due to holes in the reconstruction, making artifact detection straightforward and leading to high correlation scores for most metrics. However, scenes like treehill, stump, and flowers exhibit artifacts in the form of blurry or unnatural textures while maintaining similar color distributions to the ground truth. In these more subtle cases, Puzzle Similarity significantly outperforms FR metrics. We include extended results and artifact examples in our Supplemental.
Table 1: Pearson Correlation between Image Metrics (full-reference and no-reference (NR)) and Human Perception per Dataset
    bicycle     bonsai     counter     drjohnson     flowers     garden     kitchen     playroom     stump     train     treehill     truck
Full-Reference L1 0.199 0.325 0.384 0.280 0.078 0.630 0.533 0.556 0.141 0.314 0.114 0.314
L2 0.163 0.293 0.232 0.166 0.085 0.546 0.396 0.444 0.123 0.220 0.081 0.223
SSIM [37] 0.188 0.325 0.370 0.342 0.251 0.603 0.596 0.536 0.231 0.468 0.343 0.490
F
LIP [2]
0.268 0.403 0.452 0.366 0.194 0.719 0.684 0.652 0.239 0.319 0.142 0.402
LPIPS (vgg)) [49] 0.112 0.190 0.119 0.275 0.216 0.307 0.212 0.183 0.073 0.264 0.111 0.230
LPIPS (alex)) [49] 0.059 0.122 0.383 0.103 0.134 0.528 0.247 0.177 0.102 0.322 0.221 0.189
LPIPS (squeeze)) [49] 0.108 0.289 0.256 0.263 0.169 0.211 0.304 0.395 0.158 0.132 0.153 0.139
FovVideoVDP [19] 0.370 0.457 0.561 0.470 0.310 0.659 0.826 0.752 0.370 0.584 0.339 0.594
NR CNNIQA [12] 0.075 0.167 0.239 0.068 0.158 0.408 0.322 0.345 0.366 0.400 0.131 0.014
PIQE [26] 0.265 0.266 0.254 0.116 0.586 0.443 0.490 0.151 0.527 0.202 0.375 0.100
PuzzleSim (ours) 0.584 0.527 0.567 0.483 0.590 0.637 0.745 0.662 0.460 0.597 0.702 0.555
Table 2: Spearman Correlation between Image Metrics (full-reference and no-reference (NR)) and Human Perception per Dataset.
    bicycle     bonsai     counter     drjohnson     flowers     garden     kitchen     playroom     stump     train     treehill     truck
Full-Reference L1 0.151 0.237 0.231 0.209 0.051 0.391 0.347 0.501 0.118 0.298 0.191 0.288
L2 0.155 0.249 0.234 0.218 0.054 0.402 0.353 0.513 0.122 0.296 0.192 0.292
SSIM [37] 0.160 0.120 0.252 0.413 0.244 0.540 0.273 0.496 0.279 0.468 0.402 0.435
F
LIP [2]
0.205 0.335 0.250 0.272 0.203 0.466 0.438 0.564 0.210 0.294 0.216 0.365
LPIPS (vgg)) [49] 0.093 0.097 0.081 0.281 0.207 0.155 0.138 0.159 0.051 0.236 0.099 0.171
LPIPS (alex)) [49] 0.027 0.093 0.169 0.091 0.106 0.238 0.168 0.094 0.048 0.240 0.233 0.132
LPIPS (squeeze)) [49] 0.057 0.180 0.158 0.233 0.137 0.065 0.162 0.250 0.167 0.101 0.133 0.064
FovVideoVDP [19] 0.342 0.312 0.454 0.422 0.295 0.534 0.646 0.735 0.382 0.555 0.364 0.475
NR CNNIQA [12] 0.085 0.205 0.221 0.163 0.131 0.254 0.373 0.379 0.316 0.397 0.221 0.095
PIQE [26] 0.208 0.409 0.292 0.182 0.459 0.229 0.620 0.195 0.376 0.225 0.277 0.173
PuzzleSim (ours) 0.468 0.392 0.382 0.501 0.428 0.428 0.658 0.603 0.306 0.541 0.548 0.441
4.3 Comparison with Visual Difference Predictors
The explicit model of low-level human vision in VDP models usually produces strong correlations between the metric and human assessment. We compare against the state-of-the-art FovVideoVDP [19]; while the metric can take into account higher-order perceptual cues like motion and eccentricity, we disable them as we are 1) analyzing single static frames and 2) not assuming specific viewing conditions such as display size or distance to the screen. Puzzle Similarity not only matches but often surpasses FovVideoVDP, particularly in texture-rich scenes like treehill, stump, and flowers, again while relinquishing direct references.
4.4 Comparison with No-Reference Metrics
Closer to our work, we can compare with other no-reference quality metrics capable of producing spatial quality maps [45, 12]. Puzzle Similarity demonstrates superior accuracy in artifact localization, as shown by the correlation values in Table 3. While PIQE performed well in certain scenes, its overall correlation with human perception was generally lower in others, reflecting a less robust alignment with perceived artifact localization. However, while our method relies on a small subset of images from a similar distribution to the target image (e.g., the training dataset on novel view synthesis of a specific scene), metrics like PIQE do not require any extra images and attempt to generalize to any input. Still, in the context of 3D reconstruction, our metric performs better than all tested no-reference metrics.
Table 3: Aggregated correlation between Image Metrics and Human Perception with mean and standard deviation across all datasets.
Metric Pearson $\uparrow$ Spearman $\uparrow$
Full-Reference L1 $0.322_{\pm 0.203}$ $0.251_{\pm 0.157}$
L2 $0.248_{\pm 0.175}$ $0.257_{\pm 0.158}$
SSIM [37] $0.395_{\pm 0.166}$ $0.340_{\pm 0.168}$
F
LIP [2]
$0.403_{\pm 0.216}$ $0.318_{\pm 0.150}$
LPIPS (vgg) [49] $0.191_{\pm 0.147}$ $0.147_{\pm 0.114}$
LPIPS (alex) [49] $0.216_{\pm 0.169}$ $0.137_{\pm 0.107}$
LPIPS (squeeze) [49] $0.215_{\pm 0.148}$ $0.142_{\pm 0.097}$
FovVideoVDP [19] $0.524_{\pm 0.191}$ $0.460_{\pm 0.165}$
NR CNNIQA [12] $0.224_{\pm 0.174}$ $0.237_{\pm 0.153}$
PIQE [26] $0.314_{\pm 0.188}$ $0.304_{\pm 0.167}$
PuzzleSim (ours) $0.592_{\pm 0.119}$ $0.475_{\pm 0.137}$
5 Application: Progressive Automatic Artifact Inpainting
Finally, we will showcase a possible application of our metric in automatic restoration of novel images from a reconstructed scene.
Whenever it is possible to establish a visual distribution (e.g. we have a training dataset available), we can recursively use our metric to automatically identify visual outliers in novel views and remove them through inpainting.
Figure 5: Example showcase of our iterative inpainting application to enhance new views that lack ground-truth correspondences.
Our Framework
We can take a new image $\mathcal{I}$ and employ our PuzzleSim metric to obtain the quality map $\mathcal{Q}$.
[…]
To apply neural inpainting, we first need to create a binary mask from the quality map $\mathcal{Q}$, indicating the areas to be inpainted. This involves finding an optimal threshold $\tau$ that clearly distinguishes artifact regions. The effectiveness of inpainting depends on setting this mask carefully. If the mask is too large, the inpainting may inadvertently remove clean parts of the scene. If the mask is too small, artifacts might be left untouched. In order to automatically find a balanced threshold, we use a conservative, iterative approach to refine the test image based on the assumption that artifacts have below-average quality scores. For an initial threshold, we select $N=50$ candidate values, uniformly spaced between the lowest and mean quality scores that we use to threshold the quality map.
$\tau_{i}=$ $\min(\mathcal{Q})+\frac{i}{N-1}(\text{mean}(\mathcal{Q})-\min(\mathcal{Q}))$ (7)
$M^{(h,w)}_{i}=$$\begin{cases}1&\text{if }\mathcal{Q}^{(h,w)}\leq\tau_{i}\\ 0&\text{if }\mathcal{Q}^{(h,w)}>\tau_{i}\end{cases}$
with $i$, $h$, and $w$ representing indices where $i=0,\dots,N-1$, $h=1,\dots,H_{\mathcal{I}}$, and $w=1,\dots,W_{\mathcal{I}}$. We generate inpaintings using all $N$ masks and recompute PuzzleSim for each option. The quality of each inpainted candidate is evaluated by calculating the average quality difference within the inpainted region, denoted as $\delta_{i}$. To discourage overly large masks, we add a regularization term that penalizes them. Further details on the mathematical definitions of $\delta_{i}$ and the regularization term are provided in the Supplementary material.
We then select the candidate that maximizes $\delta$. After determining the initial threshold, we iteratively refine the inpainted image by drawing new thresholds close to the previous one. For each new threshold, we repeat the selection procedure, but with fewer candidates sampled within a narrow range around the previous threshold. The size of this interval depends on a hyperparameter $\alpha$ and the spread of quality scores. Keeping this range small ensures stable convergence and prevents excessive inpainting that could disrupt scene geometry.
If the upper limit of the interval is below the minimum quality value $\min_{h,w}\hat{\mathcal{Q}}^{(h,w)}$, we revert to the initial sampling method in Eq. 7 since an empty mask would be meaningless and cause division by zero when computing $\delta_{i}$.
Finally, we terminate the process if no further improvement is achieved (i.e., $\max_{i}\delta_{i}\leq 0$), returning the final inpainting result. This framework guarantees a monotonic improvement in PuzzleSim quality.
In Figure 5, we showcase several novel views from the reconstructed scenes garden, kitchen, and flowers using only a fraction of the original training views (20-30%). We process this artifact-ridden new view through the iterative inpainting framework presented above. Our method successfully detects and inpaints artifacts in the original reconstruction, producing high-quality inpainting consistent with the distribution of the original scenes.
6 Limitations and Future Work
While our method demonstrates promising results, there are several limitations to consider. Finding the maximum similarity with many other vectors becomes expensive as the number of reference images and image resolution rises. Performing approximate maximum search or fitting Gaussian mixture models in the embedding space can improve computational performance [48, 11]. Furthermore, our metric is empirically calibrated, but choosing the weights to combine layers and weighting in the channel dimension in a data-driven manner could advance the metric further. The resolution at which our metric can be utilized is currently limited by the CNN backbone’s generalizability to higher resolutions. Although it is differentiable, it is unlikely to produce valuable gradients due to the max operation across many vectors. The next step would be to explore softmax alternatives to make the metric more suitable for gradient-based optimization.
7 Conclusion
In this work, we have introduced Puzzle Similarity, a no-reference visual quality metric designed to detect and localize artifacts in novel views generated by 3D scene reconstruction methods. By leveraging learned patch statistics from input views, our method can generate spatial artifact maps without needing ground-truth references, which is a significant advantage for evaluating reconstructed scenes. Furthermore, we have provided a novel dataset of human-assessed quality and artifact detection specifically tailored for 3D scene reconstruction approaches.
Our evaluation demonstrates that Puzzle Similarity outperforms traditional full-reference metrics, such as SSIM and L2 norms, as well as spatial quality no-reference metrics, like PIQE, in capturing artifacts that align with human perception. Compared to sophisticated visual difference predictors like FovVideoVDP, which leverage explicit fitted models of low-level human vision, Puzzle Similarity achieves comparable or superior performance in complex texture-rich scenes, proving its robustness across a range of artifact types and scenarios, while relinquishing the necessity of direct references.
Additionally, We demonstrate our metric applied to the problem of automatic artifact inpainting, highlighting its potential for enhancing quality in scene reconstructions. Overall, Puzzle Similarity offers an effective solution for artifact localization in 3D reconstruction, paving the way for more perceptually aligned, reference-free quality assessment in computer vision and graphics, and we are hopeful it will power many more downstream applications in the future, such as few-shot reconstruction and guided acquisition.
Acknowledgements
We would like to thank Krzysztof Wolski for making their image segmentation tool available to us, Volodymyr Kyrylov for providing the idea and first prototype of the memory-efficient implementation, and Sophie Kergaßner for designing figures. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement N°804226 PERDY), from the Swiss National Science Foundation (SNSF, Grant 200502) and an academic gift from Meta.